A and B having equal skill are playing a game of best of 5 points. After A has won 2 points and B has won 1 point, what is the probability that A will win the game?

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The required probabilty will be 2/5

subrat
6 years ago
The answer is 3/4 and not 2/5. Let`s see how! Total possible cases are A wins the next point and wins the game, B wins the next two point and wins the game, B wins the next point followed by A winning the next point. As for the 4th point, as both players have equal skill, therefore probability for the next point is 1/2. When A wins the next point P=1/2, when B wins the next point followed by A winning the next point is P=(1/2)*(1/2)=1/4. Net probability by which A wins the game is hence 3/4.